Prove that n•P(n−1,n−1)= P(n,n).


1
Expert's answer
2021-10-15T09:30:52-0400

Solution:

"LHS=n.P(n-1,n-1)
\\=n\times \dfrac{(n-1)!}{(n-1-n+1)!}
\\=n\times \dfrac{(n-1)!}{(0)!}
\\=n\times \dfrac{(n-1)!}{1}
\\=n(n-1)!
\\=n!
\\=\dfrac{n!}{0!}
\\=\dfrac{n!}{(n-n)!}
\\=P(n,n)
\\=RHS"

Hence, proved.